The third fundamental form of a surface σ(u, v) is
||N̂u|| ^2 du^2 + 2N̂u.N̂v dudv + ||N̂v||^2 dv^2
where N̂ (u, v) is the standard unit normal to σ(u, v). Let FIII be the associated 2 × 2
symmetric matrix.
Show that FIII = FIIF^−1I FII , where FI and FII are the 2 × 2 symmetric matrices
associated with the first and the second fundamental forms, respectively
1
Expert's answer
2020-12-07T20:43:21-0500
"Solution"

From the introductions to First and second fundamental forms (https://warwick.ac.uk/fac/sci/maths/people/staff/weiyi_zhang/ma3d92018fall.pdf)

We prefer "E,F, and\ G" to the first fundamental form and "L,M,N" to the second fundamental form, hence;


"I=Edu^2+ 2Fdudv+Gdv^2,\ II=Ldu^2+ 2Mdudv+Ndv^2."

We denote


"\digamma_{I}=\begin{pmatrix}
 E & F \\
 F & G
\end{pmatrix}, 
\digamma_{II}=\begin{pmatrix}
 L & M \\
 M & N
\end{pmatrix}"

From the above concept, we deduced the first, second and third fundamental forms replacing "Edu^2+ 2Fdudv+Gdv^2" with "a_{11}(du^1)^2+2a_{12}du^1du^2+a_{22}(du^2)^2" , "II=Ldu^2+ 2Mdudv+Ndv^2" with "b_{11}(du^1)^2+2b_{12}du^1du^2+b_{22}(du^2)^2" and used "d\overline{n}^2=(\overline{n_{u^1}})^2(du^1)^2+2\overline{n_{u^1}}\overline{n_{u^2}}du^1du^2+(\overline{n_{u^2}})^2(du^2)^2" for the third fundamental form



"I = dS^2=a_{11}(du^1)^2+2a_{12}du^1du^2+a_{22}(du^2)^2\\
a_{ij}=(\overline{r_i}, \overline{r_j}); i,j=1,2\\
II=b_{11}(du^1)^2+2b_{12}du^1du^2+b_{22}(du^2)^2\\
b_{ij}=(\overline{r_{ij}},\overline{n}); i,j=1,2\\
III=d\overline{n}^2=(\overline{n_{u^1}})^2(du^1)^2+2\overline{n_{u^1}}\overline{n_{u^2}}du^1du^2+(\overline{n_{u^2}})^2(du^2)^2\\
\begin{pmatrix}
 (\overline{n_{u^1}})^2 & \overline{n_{u^1}}\overline{n_{u^2}} \\
 \overline{n_{u^1}}\overline{n_{u^2}} & (\overline{n_{u^2}})^2 
\end{pmatrix}
= 
\begin{pmatrix}
 b_{11} & b_{12} \\
 b_{12} & b_{22} 
\end{pmatrix}
\cdot
\begin{pmatrix}
 a_{11} & a_{12} \\
 a_{12} & a_{22}
\end{pmatrix}^{-1}
\cdot
\begin{pmatrix}
 b_{11} & b_{12} \\
 b_{12} & b_{22} 
\end{pmatrix}\\
A=\begin{pmatrix}
 a_{11} & a_{12} \\
 a_{12} & a_{22}
\end{pmatrix}\\
A^{-1}=\begin{pmatrix}
 \frac{a_{22}}{\Delta} & -\frac{a_{12}}{\Delta} \\
 -\frac{a_{12}}{\Delta} & \frac{a_{11}}{\Delta}
\end{pmatrix}\\
\Delta=|A|=a_{11}a_{22}-a_{12}^2"